# Size-consistent implementation of Hamiltonian simulation-based quantum-selected configuration interaction method for the supramolecular approach

### v1

@@ Abstract
The quantum-selected configuration interaction (QSCI) method is a promising approach for large-scale quantum chemical calculations on currently available quantum hardware. However, its naive implementation lacks size consistency, which is essential for accurate intermolecular interaction energy calculations using the supramolecular approach. Here, we present a size-consistent implementation of QSCI by sampling Slater determinants for the dimer in the localized molecular orbital basis, constructing the subspaces for the monomers and dimer, and augmenting the dimer subspace with additional determinants required for size consistency. Implemented within the Hamiltonian simulation-based QSCI (HSB-QSCI) framework, our method numerically satisfies size consistency for 4H/8H clusters, the FH dimer, and the FH–H2O system. Application to intermolecular interaction energy calculations of hydrogen-bonded FH dimer and FH–H2O demonstrates that our approach reproduces complete active space-configuration interaction (CAS-CI) values with errors below 0.04 kcal mol-1.
@@ I Introduction
Quantum computation holds the potential to revolutionize our society. Among its diverse applications, sophisticated quantum chemical calculations are considered one of the most promising. Accurate quantum chemical calculations can accelerate drug and materials discovery, elucidate reaction mechanisms in enzymes and biomolecules, and so on. Quantum phase estimation (QPE) [1, 2, 3] and variational quantum eigensolver (VQE) [4] are two major quantum computational approaches, targeting the fault-tolerant quantum computing (FTQC) and noisy intermediate-scale quantum (NISQ) eras, respectively, and both methods have been extensively investigated from theoretical and hardware demonstration perspectives. However, QPE typically requires quantum circuit that are prohibitively deep for even the most advanced quantum hardware available today, rendering its application to industrially relevant molecules extremely challenging until FTQC becomes a reality. On the other hand, VQE can be executed on NISQ devices, but it faces several significant challenges, such as substantial measurement overhead [5], the presence of barren plateaus [6], and so on.
@@
The quantum-selected configuration interaction (QSCI) method was proposed to perform large-scale quantum chemical calculations using currently available quantum hardware [7]. In QSCI, a quantum computer is used to sample Slater determinants that make significant contributions to the wave function of the target electronic state. The information from these sampled Slater determinants is then transferred to a classical computer, where the subspace Hamiltonian is constructed and diagonalized to obtain the energy and wave function. The key step in QSCI is the preparation of a quantum state suitable for sampling important Slater determinants. Initially, quantum state preparation methods based on VQE and adiabatic state preparation were proposed for this purpose [7]. In 2025, Robledo-Moreno and coworkers introduced a method for quantum state preparation based on the local unitary cluster Jastrow (LUCJ) ansatz [8] with classically computed coupled cluster singles and doubles (CCSD) excitation amplitudes. They demonstrated QSCI calculations using up to 77 qubits on superconducting quantum processors, aided by an error mitigation technique called self-consistent configuration recovery [9]. Recently, we proposed a quantum state preparation method for QSCI based on Hamiltonian simulation (HSB-QSCI) [10], in which important Slater determinants are sampled from quantum states generated by real-time evolution of an approximate wave function. HSB-QSCI offers several advantages over the original QSCI: (1) a simple wave function such as a Hartree–Fock (HF) wave function can be used as the initial state for time evolution, (2) VQE variational optimization is not required, (3) different samples can be obtained by varying the evolution time, (4) long-time evolution is challenging to simulate on classical computers, and (5) higher-order excited configurations can be sampled even with shorter evolution times. We have also reported proof-of-principle demonstrations of HSB-QSCI for one-dimensional carbon chain molecules (carbynes), utilizing up to 36 qubits of an IBM superconducting quantum processor, with the aid of matrix product operator (MPO)-based classical optimization of the quantum circuit for the time evolution operator [11]. Shortly after our publication, methods based on the same concept were reported by other groups [12, 13, 14]. Notably, Weaving and coworkers reported that time evolution-based state preparation for QSCI yields a more compact wave function expansion than conventional selected CI in SiH4 system [15].
@@
Although QSCI has been actively investigated from both theoretical and experimental perspectives [16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32], it suffers from a serious drawback: QSCI is not size consistent unless special care is taken. In its naive implementation, the energy of a dimer composed of non-interacting monomers is generally not equal to the sum of the energies of the individual monomers. In fact, as discussed in the Supporting Information, the number of quantum circuit executions must increase exponentially with system size even approximately satisfy the size consistency condition. Consequently, the straightforward application of QSCI to the calculation of intermolecular interaction energies using the supramolecular approach is fundamentally impractical. In the supramolecular approach, the intermolecular interaction energy $E_{\text{int}}$ is calculated using the following equation:
@@
Here, $E_{\text{A}\cdots\text{B}}$ denotes the energy of the dimer composed of monomers A and B at a sufficiently large inter-monomer distance. Hereafter, we refer to the approach based on equation (2) as the “dimer approach”. In 2025, Kaliakin and coworkers reported QSCI-based calculations of intermolecular interaction energies for water and methane dimers using the dimer approach [19]. However, it is evident that the computational cost is lower for the supramolecular approach than for the dimer approach, since the size of the subspace Hamiltonian to be diagonalized is smaller for monomers than for the dimer. It is also important to note that if Slater determinants are sampled independently for different geometries, the electronic configurations included in the subspace Hamiltonian may differ between geometries, potentially resulting in discontinuities in the potential energy curve. In this work, we propose a size-consistent implementation of HSB-QSCI (sc-HSB-QSCI) to overcome these challenges.
II Theory
@@
In conventional HSB-QSCI [10], Slater determinants used for subspace Hamiltonian diagonalization are sampled from quantum states generated by real-time evolution of an approximate wave function. To expand the subspace, measurement outcomes obtained at different evolution times, $k\Delta t$ for $k\leq K$, are combined. For example, HSB-QSCI with $K=3$ indicates that the subspace Hamiltonian is constructed using the measurement results from quantum circuit executions at $k$ = 1, 2, and 3.
@@
A schematic illustration of subspace construction in sc-HSB-QSCI is shown in Figure 1. Note that the following subspace construction approach is not limited to HSB-QSCI and can be universally applied to any selected CI framework. The strategy of sc-HSB-QSCI is as follows: (1) Sample Slater determinants for the dimer in the localized molecular orbital basis. (2) For each sampled bit string, check the electronic configuration of the dimer. If the configuration involves only intra-monomer excitations, split the bit string into those corresponding to each monomer and add them to the respective monomer subspace in the localized molecular orbital basis. If inter-monomer charge-transfer excitations are present, add the bit string to the dimer subspace without splitting. (3) Perform subspace Hamiltonian diagonalization for the monomers using the Slater determinants generated in the previous step. (4) The dimer subspace Hamiltonian is constructed by including all possible combinations of the electronic configurations from monomer 1 and monomer 2 subspaces, as well as the Slater determinants corresponding to charge-transfer excitations. With this approach, electronic configurations required for size consistency but not sampled are systematically added, thereby ensuring that size consistency is automatically satisfied.
III Computational conditions
As a proof-of-principle demonstration of sc-HSB-QSCI, we performed numerical simulations on a classical computer. The target systems included a dimer of square tetrahydrogen (4H) clusters and hydrogen-bonded systems (FH dimer and FH–H2O). The H–H bond length in the 4H cluster was set to 1.0583 Å (= 2.0 Bohr), and the 8H cluster was constructed by placing two 4H clusters in a cuboid arrangement with an inter-monomer distance of 100 Å. Note that this system has previously been used to investigate size consistency in VQE with a unitary coupled cluster singles and doubles (UCCSD) ansatz [33], and in QPE [34]. The localized molecular orbitals of the 8H cluster were manually generated by taking appropriate linear combinations of RHF/STO-3G [35] canonical orbitals. For sc-HSB-QSCI, all molecular orbitals were included in the Hamiltonian simulation, and the interaction energies were compared with those from full-configuration interaction (full-CI) calculations.
@@ III Computational conditions
For the hydrogen-bonded systems, geometry optimizations of the dimers were performed at the PBE0 [36]/aug-cc-pVDZ [37, 38] level of theory. No imaginary vibrational frequencies were obtained at the optimized structure. The optimized geometries are shown in Figure 2, and the corresponding Cartesian coordinates are provided in Tables S1 and S2 in the Supporting Information. For monomer calculations, we used the optimized geometry for the dimer without further performing geometry optimization of the monomers. To assess size consistency, we also generated geometries for the dimers without inter-monomer interaction by increasing the inter-monomer distance to 100 Å. Localized molecular orbitals for both dimers and monomers were generated using the Pipek–Mezey method [39] at the RHF/aug-cc-pVDZ level. In the sc-HSB-QSCI simulations, only valence orbitals were considered, with active space sizes of (16e, 10o) for the FH dimer, (8e, 5o) for the FH monomer, (16e, 11o) for the FH–H2O, and (8e, 6o) for the H2O. Here, (NNe, MMo) indicates an active space consisting of $N$ electrons and $M$ molecular orbitals. The active orbitals for the monomers and dimers of the FH dimer and FH–H2O are shown in Figure S2 in Supporting Information. All DFT and RHF calculations were performed using the GAMESS-US program package [40].
@@
Numerical simulations for sc-HSB-QSCI were performed using our in-house Python3 code, developed with the OpenFermion [41], Cirq [42], and PyCI [43] libraries. Details of the HSB-QSCI implementation can be found in Reference [10]. The Jordan–Wigner transformation [44] was used for fermion–qubit encoding. The RHF wave function served as the initial state for time evolution, and the quantum circuit for the time evolution operator $U=e^{-iH\Delta t}$ with $\Delta t=1$ was constructed using the second-order Trotter–Suzuki decomposition with a single Trotter step. Unless otherwise noted, the number of shots for quantum circuit execution was set to $1\times 10^{4}$, and $K$ is varied from 1 to 5 and 10. In sc-HSB-QSCI, symmetry completion [10] was performed prior to constructing the subspace Hamiltonian matrix. In this step, any open-shell Slater determinants required to span the eigenfunctions of the electron spin ${\bf S}^{2}$ operator were added if missing.
@@
For the hydrogen-bonded FH dimer and FH–H2O systems, we also calculated intermolecular interaction energies using HCI in conjunction with the propose approach. The HCI calculations were performed using the PyCI [43] library.
IV Results and Discussion
IV.1 Size consistency of sc-HSB-QSCI

### v2

@@ Abstract
The quantum-selected configuration interaction (QSCI) method is a promising approach for large-scale quantum chemical calculations on currently available quantum hardware. However, its naive implementation lacks size consistency, which is essential for accurate intermolecular interaction energy calculations using the supramolecular approach. Here, we present a size-consistent implementation of QSCI by sampling Slater determinants for the dimer in the localized molecular orbital basis, constructing the subspaces for the monomers and dimer, and augmenting the dimer subspace with additional determinants required for size consistency. Implemented within the Hamiltonian simulation-based QSCI (HSB-QSCI) framework, our method numerically satisfies size consistency for 4H/8H/12H clusters, the FH dimer, and the FH–H2O system. Application to intermolecular interaction energy calculations of hydrogen-bonded FH dimer and FH–H2O demonstrates that our approach reproduces complete active space-configuration interaction (CAS-CI) values with errors below 0.04 kcal mol-1.
@@ I Introduction
Quantum computation holds the potential to revolutionize our society. Among its diverse applications, sophisticated quantum chemical calculations are considered one of the most promising. Accurate quantum chemical calculations can accelerate drug and materials discovery, elucidate reaction mechanisms in enzymes and biomolecules, and so on. Quantum phase estimation (QPE) [Abrams-1999, Aspuru-Guzik-2005, Nielsen-Chuang] and variational quantum eigensolver (VQE) [Peruzzo-2014] are two major quantum computational approaches, targeting the fault-tolerant quantum computing (FTQC) and noisy intermediate-scale quantum (NISQ) eras, respectively, and both methods have been extensively investigated from theoretical and hardware demonstration perspectives. However, QPE typically requires quantum circuit that are prohibitively deep for even the most advanced quantum hardware available today, rendering its application to industrially relevant molecules extremely challenging until FTQC becomes a reality. On the other hand, VQE can be executed on NISQ devices, but it faces several significant challenges, such as substantial measurement overhead [Gonthier-2022], the presence of barren plateaus [McClean-2018], and so on.
@@
The quantum-selected configuration interaction (QSCI) method was proposed to perform large-scale quantum chemical calculations using currently available quantum hardware [Kanno-2023]. In QSCI, a quantum computer is used to sample Slater determinants that make significant contributions to the wave function of the target electronic state. The information from these sampled Slater determinants is then transferred to a classical computer, where the subspace Hamiltonian is constructed and diagonalized to obtain the energy and wave function. The key step in QSCI is the preparation of a quantum state suitable for sampling important Slater determinants. Initially, quantum state preparation methods based on VQE and adiabatic state preparation were proposed for this purpose [Kanno-2023]. In 2025, Robledo-Moreno and coworkers introduced a method for quantum state preparation based on the local unitary cluster Jastrow (LUCJ) ansatz [Motta-2023] with classically computed coupled cluster singles and doubles (CCSD) excitation amplitudes. They demonstrated QSCI calculations using up to 77 qubits on superconducting quantum processors, aided by an error mitigation technique called self-consistent configuration recovery [Robledo-Moreno-2025]. Recently, we proposed a quantum state preparation method for QSCI based on Hamiltonian simulation (HSB-QSCI) [Sugisaki-2025], in which important Slater determinants are sampled from quantum states generated by real-time evolution of an approximate wave function. HSB-QSCI offers several advantages over the original QSCI: (1) a simple wave function such as a Hartree–Fock (HF) wave function can be used as the initial state for time evolution, (2) VQE variational optimization is not required, (3) different samples can be obtained by varying the evolution time, (4) long-time evolution is challenging to simulate on classical computers, and (5) higher-order excited configurations can be sampled even with shorter evolution times. We have also reported proof-of-principle demonstrations of HSB-QSCI for one-dimensional carbon chain molecules (carbynes), utilizing up to 36 qubits of an IBM superconducting quantum processor, with the aid of matrix product operator (MPO)-based classical optimization of the quantum circuit for the time evolution operator [Kanno-2025]. Shortly after our publication, methods based on the same concept were reported by other groups [Mikkelsen-2024, Yu-2025, Piccinelli-2025]. Notably, Weaving and coworkers reported that time evolution-based state preparation for QSCI yields a more compact wave function expansion than conventional selected CI in SiH4 system [Weaving-2025].
@@
Although QSCI has been actively investigated from both theoretical and experimental perspectives [Nakagawa-2024, Liepuoniute-2025, Nutzel-2025, Kaliakin-2024, Barison-2025, Shajan-2025, Reinholdt-2025, Kaliakin-2025, Bauer-2025, Duriez-2025, Ohgoe-2025, Nogaki-2025, Bickley-2025, Chen-2025, Shivpuje-2025, Smith-2025, Matsuyama-2025], it suffers from a serious drawback: QSCI is not size consistent unless special care is taken. In its naive implementation, the energy of a dimer composed of non-interacting monomers is generally not equal to the sum of the energies of the individual monomers. In fact, as discussed in the Supporting Information, the number of quantum circuit executions must increase exponentially with system size even approximately satisfy the size consistency condition. Consequently, the straightforward application of QSCI to the calculation of intermolecular interaction energies using the supramolecular approach is fundamentally impractical. In the supramolecular approach, the intermolecular interaction energy $E_{\text{int}}$ is calculated using the following equation:
@@
Here, $E_{\text{A}\cdots\text{B}}$ denotes the energy of the dimer composed of monomers A and B at a sufficiently large inter-monomer distance. Hereafter, we refer to the approach based on equation (2) as the “dimer approach”. In 2025, Kaliakin and coworkers reported QSCI-based calculations of intermolecular interaction energies for water and methane dimers using the dimer approach [Kaliakin-2024]. However, it is evident that the computational cost is lower for the supramolecular approach than for the dimer approach, since the size of the subspace Hamiltonian to be diagonalized is smaller for monomers than for the dimer. It is also important to note that if Slater determinants are sampled independently for different geometries, the electronic configurations included in the subspace Hamiltonian may differ between geometries, potentially resulting in discontinuities in the potential energy curve. It should be emphasized that these issues also arise in other QSCI variants, as long as the selected CI framework is employed. In this work, we propose a size-consistent implementation of HSB-QSCI (sc-HSB-QSCI) to overcome these challenges.
The rest of the paper is organized as follows. Section II provides a brief overview of quantum computational approaches to intermolecular interaction energy calculations reported to date, highlighting both their potential advantages and methodological challenges. In Section III, we discuss the theoretical foundations of the size-consistent implementation of HSB-QSCI. Section IV describes the computational conditions for proof-of-principle numerical simulations performed using a statevector simulator. Section V presents the results of these simulations, including validation of size consistency and application of the proposed method to intermolecular interaction energy calculations for hydrogen-bonded systems. Section VI concludes with a summary of our findings and outlines future prospects.
II Quantum computational approaches for intermolecular interaction energy calculations
In this section, we present a brief overview of quantum computational approaches to intermolecular interaction energy calculations.
There are two major approaches to intermolecular interaction energy calculations, namely the symmetry-adapted perturbation theory (SAPT) and the supramolecular approach. SAPT is a perturbation theory-based method that evaluates intermolecular interaction energies using the wave functions of the individual monomers as the unperturbed states. Because SAPT defines the wave functions and energies for each monomer and introduces interaction terms perturbatively, the intermolecular interaction energy becomes zero when the molecules are infinitely separated, thereby ensuring size consistency automatically. Within the VQE framework, one- and two-electron reduced density matrices (1RDM and 2RDM, respectively) can be obtained via sampling on a quantum computer, making the application of VQE to SAPT relatively straightforward [Malone-2022, Loipersberger-2023]. In the context of QPE, Cortes and coworkers have reported an attempt to combine QPE with SAPT [Cortes-2024]. However, since SAPT requires the calculation of expectation values of the electrostatic ($\hat{V}$), exchange ($\hat{P}$), and electrostatic exchange ($\widehat{VP}_{S}$) operators, it must be integrated with expectation value estimation algorithms [Steudtner-2023]. This makes practical implementation on quantum hardware challenging, unless fault-tolerant quantum computing (FTQC) is realized. To the best of our knowledge, there are no reports of studies combining the QSCI method with SAPT.
In the supramolecular approach, the intermolecular interaction energy is calculated using equation (1), as discussed in the Introduction section. While this approach appears straightforward for evaluating intermolecular interaction energies, particular attention must be paid in quantum computations to ensure that the size consistency condition is satisfied. In fact, although the unitary coupled cluster (UCC) ansatz in VQE has generally been regarded as size-consistent due to its foundation in coupled cluster theory, our recent work [Sugisaki-2024a] revealed that size consistency in VQE-UCC can be broken when the Trotter decomposition is used to construct the quantum circuit. We found that using molecular orbitals localized on each monomer is essential for maintaining size consistency. In the case of QPE, when the quantum circuit for the time evolution operator is constructed via Trotter decomposition, not only the order of the Trotter decomposition (first or second order) and the number of Trotter slices, but also the choice of molecular orbitals for wave function expansion (RHF canonical orbitals or localized orbitals) and the term ordering in the Trotterized time evolution operator can impact size consistency condition [Sugisaki-2024b]. Intermolecular interaction energy calculations based on the supramolecular approach using QPE have been reported recently [Tachi-2025].
Thus, based on an overview of the field of quantum computational chemistry, the calculation of intermolecular interaction energies using quantum computers remains an area of active development. Consequently, the development of computational methods capable of accurately and efficiently evaluating intermolecular interaction energies on quantum computers represents a significant and ongoing challenge.
III Theory
@@
In conventional HSB-QSCI [Sugisaki-2025], Slater determinants used for subspace Hamiltonian diagonalization are sampled from quantum states generated by real-time evolution of an approximate wave function. To expand the subspace, measurement outcomes obtained at different evolution times, $k\Delta t$ for $k\leq K$, are combined. For example, HSB-QSCI with $K=3$ indicates that the subspace Hamiltonian is constructed using the measurement results from quantum circuit executions at $k$ = 1, 2, and 3. In the original proposal of the QSCI method [Kanno-2023], only the $R$ most frequently observed Slater determinants are selected to define the subspace. In our HSB-QSCI implementation [Sugisaki-2025], we do not impose such a threshold, and all sampled Slater determinants are included in the subspace.
Because the wave functions are simultaneous eigenfunctions of both the Hamiltonian and the electron spin ${\bf S}^{2}$ operator (hereafter referred to as spin eigenfunctions), whereas open-shell Slater determinants with spin-$\beta$ unpaired electrons are not. Without loss of generality, we assume $n_{\alpha}\geq n_{\beta}$, where $n_{\alpha}$ and $n_{\beta}$ denote the numbers of spin-$\alpha$ and spin-$\beta$ electrons, respectively. Due to the inherently stochastic nature of sampling on a quantum computer, it is possible that only a subset of the Slater determinants required to construct spin eigenfunctions will be sampled. To address this, HSB-QSCI introduces a symmetry completion step prior to constructing the subspace Hamiltonian on a classical computer. During this step, if the $|2\alpha\beta\alpha\beta 0\rangle$ configuration (where 2, $\alpha$, $\beta$, and 0 indicate that the corresponding molecular orbital is doubly occupied, singly occupied by a spin-$\alpha$ electron, singly occupied by a spin-$\beta$ electron, and unoccupied, respectively) is sampled, then the configurations $|2\alpha\alpha\beta\beta 0\rangle$, $|2\alpha\beta\beta\alpha 0\rangle$, $|2\beta\alpha\alpha\beta 0\rangle$, $|2\beta\beta\alpha\alpha 0\rangle$, and $|2\beta\alpha\beta\alpha 0\rangle$ are also added if they were not sampled.
@@
A schematic illustration of subspace construction in sc-HSB-QSCI is shown in Figure 1. Note that the following subspace construction approach is not limited to HSB-QSCI and can be universally applied to any selected CI framework. The strategy of sc-HSB-QSCI is as follows: (1) Sample Slater determinants for the dimer in the localized molecular orbital basis. (2) For each sampled bit string, check the electronic configuration of the dimer. If the configuration involves only intra-monomer excitations from the initial state, split the bit string into those corresponding to each monomer and add them to the respective monomer subspace in the localized molecular orbital basis. If inter-monomer charge-transfer excitations are present, add the bit string to the dimer subspace without splitting. Note that when the Jordan–Wigner transformation is used for fermion–qubit encoding, the Hamming weight corresponds to the number of electrons. As a result, it becomes straightforward to assign electronic configurations to intra-monomer or inter-monomer excitations. (3) Perform subspace Hamiltonian diagonalization for the monomers using the Slater determinants generated in the previous step. (4) The dimer subspace Hamiltonian is constructed by including all possible combinations of the electronic configurations from monomer 1 and monomer 2 subspaces, as well as the Slater determinants corresponding to charge-transfer excitations. With this approach, electronic configurations required for size consistency but not sampled are systematically added, thereby ensuring that size consistency is automatically satisfied.
In sc-HSB-QSCI, it is assumed that the molecular orbitals of the monomer and dimer are nearly identical. To satisfy this condition, we applied molecular orbital localization to both the monomers and the dimer in step 1. However, when two monomers are in close proximity and the inter-monomer interaction is significant, as in the hydrogen-bonded systems studied here, the localized orbitals cannot be completely confined to each monomer and instead exhibit small amplitudes on the other monomer, which may affect size consistency. To improve the locality of the molecular orbitals on each monomer, introducing a penalty term to the cost function when molecular orbitals are delocalized over both monomers could be a viable approach, which remains a topic for future investigation.
IV Computational conditions
As a proof-of-principle demonstration of sc-HSB-QSCI, we performed numerical simulations on a classical computer. The target systems included a dimer and a trimer of square tetrahydrogen (4H) clusters, hydrogen-bonded systems (FH dimer and FH–H2O). The H–H bond length in the 4H cluster was set to 2.0 Bohr ($\approx$ 1.0583 Å), and the dimer (8H cluster) and trimer (12H cluster) were constructed by placing two and three 4H clusters in a cuboid arrangement with an inter-monomer distance of 100 Å. Note that 4H/8H clusters have previously been used to investigate size consistency in VQE with a unitary coupled cluster singles and doubles (UCCSD) ansatz [Sugisaki-2024a], and in QPE [Sugisaki-2024b]. The localized molecular orbitals of the 8H and 12H clusters were manually generated by taking appropriate linear combinations of RHF/STO-3G [STO-3G] canonical orbitals. For sc-HSB-QSCI, all molecular orbitals were included in the Hamiltonian simulation, and the interaction energies were compared with those from full-configuration interaction (full-CI) calculations.
@@ IV Computational conditions
For the hydrogen-bonded systems, geometry optimizations of the dimers were performed at the PBE0 [PBE0]/aug-cc-pVDZ [cc-pVDZ, aug_for_cc-pVDZ] level of theory. No imaginary vibrational frequencies were obtained at the optimized structure. The optimized geometries are shown in Figure 2, and the corresponding Cartesian coordinates are provided in Tables S1 and S2 in the Supporting Information. For monomer calculations, we used the optimized geometry for the dimer without further performing geometry optimization of the monomers. To assess size consistency, we also generated geometries for the dimers without inter-monomer interaction by increasing the inter-monomer distance to 100 Å. Localized molecular orbitals for both dimers and monomers were generated using the Pipek–Mezey method [Pipek-1989] at the RHF/aug-cc-pVDZ level. In the sc-HSB-QSCI simulations, only valence orbitals were considered, with active space sizes of (16e, 10o) for the FH dimer, (8e, 5o) for the FH monomer, (16e, 11o) for the FH–H2O, and (8e, 6o) for the H2O. Here, (NNe, MMo) indicates an active space consisting of $N$ electrons and $M$ molecular orbitals. The active orbitals for the monomers and dimers of the hydrogen-bonded FH dimer and FH–H2O are shown in Figure S2 in Supporting Information. All DFT, RHF, and MP2 calculations were performed using the GAMESS-US program package [GAMESS].
@@
Numerical simulations for sc-HSB-QSCI were performed using our in-house Python3 code, developed with the OpenFermion [OpenFermion], Cirq [Cirq], and PyCI [PyCI] libraries. Details of the HSB-QSCI implementation can be found in Reference [Sugisaki-2025]. The Jordan–Wigner transformation [Jordan-1928] was used for fermion–qubit encoding. The RHF wave function served as the initial state for time evolution, and the quantum circuit for the time evolution operator $U=e^{-iH\Delta t}$ with $\Delta t=1$ was constructed using the second-order Trotter–Suzuki decomposition with a single Trotter step. Unless otherwise noted, the number of shots for quantum circuit execution was set to $1\times 10^{4}$, and $K$ is varied from 1 to 5 and 10. In sc-HSB-QSCI, symmetry completion [Sugisaki-2025] was performed prior to constructing the subspace Hamiltonian matrix, as discussed in the previous section.
@@
For the hydrogen-bonded FH dimer and FH–H2O systems, we also calculated intermolecular interaction energies using heat-bath configuration interaction (HCI) in conjunction with the propose approach. The HCI calculations were performed using the PyCI [PyCI] library.
V Results and Discussion
V.1 Size consistency of sc-HSB-QSCI